# Strict quantization for compact pseudo-Kähler manifolds and group actions

### v1

@@ 4 Quotients and complexification of groups
$\underline{\mathfrak{g}}={\rm Span\,}(\xi_{X};\,\xi\in\mathfrak{g})$
and
$\underline{\mathfrak{g}}^{\perp_{b}}=\left\{v\in HX;\,b(\xi_{X},v)=0,\ \ \forall\xi_{X}\in\underline{\mathfrak{g}}\right\}\,.$
By [HH, Section 2.5], we observe that when restricted to $\mathfrak{g}\times\mathfrak{g}$, the bilinear form $b$ is nondegenerate on $Y$. Furthermore,
$\pi_{Y}\left(\underline{\mathfrak{g}}_{x}\right)=T_{\pi(x)}\Phi^{-1}(0)\,$
and
$\left\{{\xi}_{M}(\pi(x))\in T_{\pi(x)}M\,:\,\xi\in\sqrt{-1}\,\mathfrak{g}\right\}$
$M_{s}=\bigcup_{g\in G^{\mathbb{C}}}g\,U$
then $M_{s}$ is itself open. Thus, the stabilizer algebra of $x$ in $\mathfrak{g}^{\mathbb{C}}$ is zero; so the action of $G^{\mathbb{C}}$ on $M_{s}$ is locally free. To show that $G^{\mathbb{C}}$ acts freely on $M_{s}$ one can proceed as in the second part of [GS, Theorem 4.5] at page 527. ∎
Now, we shall need the following theorem.
@@
The set $M\setminus M_{s}$ is contained in a complex sub-manifold of $M$ of complex codimension greater than
[…]
This theorem is an analogue of [GS, Theorem 5.7] for pseudo-Riemannian manifolds. Note that here the existence theorem [GS, Theorem 5.6] is not true in general for the pseudo-Riemannian case: we have to replace the spaces of smooth sections of $L^{k}$ with space of $L^{k}$-valued $(0,\,n_{-})$-forms, see Theorem 4.5. Then, Theorem 4.3 follows by combining Theorem 4.4 and 4.5 below.
@@
Let $s$ be a $G$-invariant $L^{k}$-valued $(0,\,n_{-})$-forms on $M$. Then, given $m\in M$ if $s(m)\neq 0$, then $m\in M_{s}$.
@@
See the proofs of Theorems 5.3 and 5.4 in [GS], they hold true in our setting up to replacing the space of $G$-invariant holomorphic sections of $L^{k}$ with the space of $G$-invariant $L^{k}$-valued $(0,\,n_{-})$-forms. ∎
Theorem 4.5.
If the set $\Phi^{-1}(0)$ is non-empty and zero is a regular value of $\Phi$, then for some $k$, there exists a global non-vanishing holomorphic $G$-invariant $L^{k}$-valued $(0,\,n_{-})$-forms.
Proof.
This is a consequence of the asymptotic expansions of $G$-invariant Szegő kernel, see [HH]. Indeed note that
$\dim_{\mathbb{C}}H^{G}_{k}(X)=\int_{X}S^{G}_{k}(x,\,x)\,\mathrm{dV}_{X}(x)\,.$
Now, since $S^{G}_{k}$ has a non-trivial asymptotic expansion on $\Phi^{-1}(0)$ in decreasing integer power of $k$, whose leading term has degree $n$. Thus, for $k$ sufficiently large $H^{G}_{k}(X)$ can not be empty. ∎
Now, “quantization commutes with reduction” follows as in [GS] by combining [GS, Theorem 5.7] (the analogous result here is Theorem 4.3) and [GS, Theorem 5.2].
@@
$\left\lVert I_{k}^{G}[f]\right\rVert_{\infty}\leq\rVert T^{G}_{k}[\widetilde{f}]\lVert\leq\lVert f\rVert_{\infty}\,.$ (15)
Recall that given two $G$ invariant $(0,\,n_{-})$ forms $s$ and $t$ in $H(X)^{G}_{k}$, we denote their inner product by $\left(s|t\right)$. Now, using the Cauchy-Schwartz inequality, we obtain that for each element $x\in\mu^{-1}(0)$
$\left\lVert I^{G}_{k}[f](x)\right\rVert^{2}_{\infty}$ $=\frac{\left|\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}T^{G}_{k}[f]\,S^{G}_{k}(\cdot,\,x)\,v\right)\right|^{2}}{\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}\,S^{G}_{k}(\cdot,\,x)\,v\right)^{2}}$
$\leq\frac{\left(T^{G}_{k}[f]\,S^{G}_{k}(\cdot,\,x)\,v\bigg{|}T^{G}_{k}[f]\,S^{G}_{k}(\cdot,\,x)\,v\right)}{\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}\,S^{G}_{k}(\cdot,\,x)\,v\right)}$
$\leq\rVert T^{G}_{k}[f]\lVert^{2}$
where the last inequality follows by the definition of the operator norm, see (6). Taking the $\sup$ both sides over $\mu^{-1}(0)$, it shows the first inequality in (15).
@@
$\frac{\lVert T^{G}_{k}[f]s\rVert^{2}}{\lVert s\rVert^{2}}$ $=\frac{\left(T^{G}_{k}[f]s\,\bigg{|}T^{G}_{k}[f]\,s\right)}{\int_{X}\langle s,\,s\rangle\,\mathrm{dV}_{X}}$ (16)
$=\frac{\lVert f\rVert^{2}_{\infty}\,\left(S^{G}_{k}[f]s\,\bigg{|}S^{G}_{k}[f]\,s\right)}{\int_{X}\langle s,\,s\rangle\,\mathrm{dV}_{X}}$
$\leq\lVert f\rVert^{2}_{\infty}\,.$
Hence, it proves the second inequality (15).
Now, we are ready to prove the theorem. Since $X$ is compact, then $\mu^{-1}(0)$ is closed and hence compact. Choose $x_{e}\in\mu^{-1}(0)$ a point with
$\lvert\widetilde{f}(x_{e})\rvert=\lVert f\rVert_{\infty}\,.$
Since the Berezin transform has as a leading term the identity, we have
$\left\lvert(I^{G}_{k}[f])(x_{e})-\widetilde{f}(x_{e})\right\rvert\leq\frac{C}{k}$
for a suitable $C>0$. Thus,
$\lVert f\rVert_{\infty}-\frac{C}{k}=\left\lvert\widetilde{f}(x_{e})\right\rvert-\frac{C}{k}\leq\left\lvert\left(I^{G}_{k}[f]\right)(x_{e})\right\rvert\leq\lVert I^{G}_{k}[f]\rVert_{\infty}\,.$
Putting (15) in this last inequality, we get part (a) of Theorem 1.1.
Now, part (b) and part (c) of Theorem 3.1 are consequence of the Theorem 2.3.
6 Proof of Corollary 1.1
Let $f$ and $g$ be two $G$ and $S^{1}$ invariant smooth function on $X_{G}$. For ease of notation we write $T^{G}_{k}[f]$ for $T^{G}_{k}[\widetilde{f}]$. It is clear by Theorem 1.1 and [GH1] that $\star$ is a well-defined associative star product. In fact, by part (b) of Theorem 1.1, we get
$\left\lVert T^{G}_{k}[{\{f,\,g\}-i(C_{1}(f,g)-C_{1}(g,f))}]\right\rVert=O\left(\frac{1}{k}\right)$
and taking the limit for $k\rightarrow+\infty$ and using part (a) we get
$\lVert\{f,\,g\}-i(C_{1}(f,g)-C_{1}(g,f))\rVert_{\infty}=0$
and thus we get part (b). Similarly, by Theorem 1.1 part (c) and part (a), we get that $C_{0}(g,h)=g\cdot h$. In a similar way the associativity property is proved.
Now, we prove the uniqueness. Let $C_{j}$ and $\tilde{C}_{j}$ be two bi-differential operators inducing a star product. Note $C_{0}=\tilde{C}_{0}$ by hypothesis. By induction assume that $C_{j}=\tilde{C}_{j}$ for each $j\leq N-2$. By the results in [GH1], we have
$T_{k}^{G}[f]\circ T_{k}^{G}[g]\sim\sum_{j=0}^{N-1}k^{-j}T_{k}^{G}[C_{j}(f,\,g)]+O(k^{-N})$
@@
$T_{k}^{G}[f]\circ T_{k}^{G}[g]\sim\sum_{j=0}^{N-1}k^{-j}T_{k}^{G}[\tilde{C}_{j}(f,\,g)]+O(k^{-N})\,,$
@@
$\left\lVert\frac{1}{k^{N-1}}\,T^{G}_{k}[{C}_{N-1}(f,\,g)-\tilde{C}_{N-1}(f,\,g)]\right\rVert\leq\frac{K}{k^{N}}\,.$
Eventually, taking the limit on both sides as $k$ goes to infinity and applying part a) of Theorem 1.1, we obtain
$C_{N-1}(f,\,g)=\tilde{C}_{N-1}(f,\,g)\,.$
@@
[…]
for $j\geq 1$ and for every $f\in C^{\infty}(X_{G})^{S^{1}}$. First, by Theorem 2.1 note that
$C_{0}(1,\,g)=C_{0}(1,\,g)=g$
for every $k$. Now, recall that
$A_{N}=\hat{R}^{l}T^{G}[f]T^{G}[g]-\sum_{j=0}^{N-1}\hat{R}^{N-j}T^{G}[C_{j}(f,\,g)]$
which is an operator of Szegő type of order zero by Theorem 2.2 and it is invariant under the circle action. Thus, its symbol is a $G$ and $S^{1}$ invariant function on $X_{G}$ which, by restriction on $\mu^{-1}(0)$, defines a $S^{1}$-invariant function on $X_{G}$. By definition, it is the next element $C_{N}(f,\,g)$ in the star-product. Now, if we put $g=1$
$A_{1}=\hat{R}T^{G}[f]T^{G}[g]-\hat{R}T^{G}[f\cdot g]=0$
and hence the symbol of $A_{1}$ vanishes. By induction the claim (17) follows.
Let us now prove that the Berezin-Toeplitz star product fulfills parity. Considering the formal parameter to be real, $\nu=\overline{\nu}$, this is equivalent to
$\overline{C_{k}(f,\,g)}=C_{k}(\overline{g},\,\overline{f})\,,$ (18)
for $k\geq 0$. Note that, as an easy consequence of the definition of Berezin-Toeplitz star product, we have
$T_{k}^{G}[f]^{\dagger}=T_{k}^{G}[\overline{f}]\,.$
The star product $\overline{g}\star\overline{f}$ is given by the asymptotic expansion of the composition
$T_{k}^{G}\left[\overline{g}\right]\cdot T_{k}^{G}\left[\overline{f}\right]$ $=T_{k}^{G}[g]^{{\dagger}}\cdot T_{k}^{G}[f]^{{\dagger}}$
$=\left(T_{k}^{G}[{f}]\cdot T_{k}^{G}[g]\right)^{{\dagger}}$
$\sim\sum_{j=0}^{+\infty}k^{-j}T_{k}^{G}\left[C_{j}(f,\,g)\right]^{{\dagger}}$
$\sim\sum_{j=0}^{+\infty}k^{-j}T_{k}^{G}\left[\overline{C_{j}(f,\,g)}\right]\,.$
Thus, the claim (18) follows.
By [GH1], see also [GH2], the trace $\mathrm{Tr}_{k}$ on $H(X)^{G}_{k}$ admits a complete asymptotic expansion in decreasing integer power of $k$:
$\mathrm{Tr}_{k}\left(T^{G}_{k}[f]\right)\sim k^{d}\left(\sum_{k=0}^{+\infty}k^{-j}\,\tau_{j}(f)\right)\,,$
with $\tau_{j}(f)\in\mathbb{C}$. It induces a $\mathbb{C}[[\nu]]$-linear map
$\mathrm{Tr}\,:\,{C}^{\infty}(X_{G})^{S^{1}}[[\nu]]\rightarrow\nu^{-n}\,\mathbb{C}[[\nu]]$
such that
@@
$\mathrm{Tr}(f\star g)=\mathrm{Tr}(g\star f)\,.$ (19)
By $\mathbb{C}[[\nu]]$-linearity, we prove (19) for $f$ and $g$ in ${C}^{\infty}(X_{G})^{S^{1}}$. Note that $f\star g-g\star f$ is given by the asymptotic expansion of
$T^{G}_{k}[f]\,T^{G}_{k}[g]-T^{G}_{k}[g]\,T^{G}_{k}[f]\,.$
Hence the trace of $f\star g-g\star f$ is given by the expansion of
$\mathrm{Tr}_{k}\left(T^{G}_{k}[f]\,T^{G}_{k}[g]-T^{G}_{k}[g]\,T^{G}_{k}[f]\right)\,.$
But for every k this vanishes and thus we get (19).
Acknowledgments: The author thanks Chin-Yu Hsiao and Herrmann Hendrik for their valuable conversations. The author is a member of GNSAGA, part of the Istituto Nazionale di Alta Matematica, and expresses gratitude to the group for its support.

### v2

@@ 5.1 Quotients and complexification of groups
$b\in S^{n-d/2}_{{\rm cl}}\bigl{(}1;D\times D,T^{*0,q}M\boxtimes\bigl{(}T^{*0,q}M\bigr{)}^{*}\bigr{)},\qquad b(x,y,k)\sim\sum_{j=0}b_{j}(x,y)k^{n-j}$
in $S^{n-d/2}_{{\rm cl}}\bigl{(}1;D\times D,T^{*0,q}M\boxtimes\bigl{(}T^{*0,q}M\bigr{)}^{*}\bigr{)}$, and
$b_{j}\in\mathcal{C}^{\infty}\bigl{(}D\times D,T^{*0,q}M\boxtimes\bigl{(}T^{*0,q}M\bigr{)}^{*}\bigr{)},\qquad b_{0}=\mathbf{b_{0}}(m)P_{\det\overline{W}^{*}}$
with $\mathbf{b_{0}}(m)>0$ and $j=0,1,\dots$, and we refer to [11, Theorem 1.8] for the properties of the phase function $\Psi\in\mathcal{C}^{\infty}(D\times D)$.
@@
If the set $\Phi^{-1}(0)$ is non-empty and zero is a regular value of $\Phi$, then for some $k$, there exists a global nonzero holomorphic $G$-invariant $L^{k}$-valued $(0,n_{-})$-forms.
Proof.
Let $S_{0},\dots,S_{d_{k}}$ be an orthonormal basis of the vector space $H^{q}\bigl{(}M,L^{\otimes k}\bigr{)}^{G}$. In the standard situation, when the line bundle is positive, we consider $q=0$. Let $m\in\Phi^{-1}(0)$; since all the components transform by the same scalar under a change of frame and since
$P^{(0),G}_{k}(m,m)=\sum_{j=0}^{d_{k}}|S_{j}(m)|^{2}_{h^{L^{\otimes k}}}=O\bigl{(}k^{n-d/2}\bigr{)},$
it is easy to see that the sections $S_{j}$ do not share common zeros.
In the pseudo-Riemannian setting, we assume $q=n_{-}$. Let $m\in\Phi^{-1}(0)$, in view of Theorem 5.8, we have
$P_{k}^{(q)}(m,m)=\mathbf{b_{0}}(m)P_{\det\overline{W}^{*}}k^{n-d/2}+O\bigl{(}k^{n-d/2-1}\bigr{)},$ (5.2)
with $\mathbf{b_{0}}(m)>0$. By (5.2), it follows that the sections $S_{j}$ do not share common zeros. Since $M$ is compact, there exists $k_{0}$ such that $P^{(n_{-}),G}_{k}(m,m)$ for all $k\geq k_{0}$ and $m\in M$. Thus, by equation (5.2), there do not exist $m\in M$ such that for each $s$ in $\mathcal{Q}_{k}(M)^{G}$, $s(m)=0$ for all ${k>0}$. In particular, since $\Phi^{-1}(0)$ is non-empty, then for some $k$, there exists a global non-zero holomorphic $G$-invariant $L^{k}$-valued $(0,n_{-})$-forms. ∎
“Quantization commutes with reduction” follows as in [9] by combining [9, Theorem 5.7] (the analogous result here is Theorem 5.6) and the following analogue of [9, Theorem 5.8].
@@
Let $s$ be a vector in $\mathcal{Q}_{k}(M)^{G}$. Then $\langle s,s\rangle$ is bounded and takes its maximum value on $\Phi^{-1}(0)$.
@@
Let $m$ be a point of $M_{s}$. Then $m=gm_{0}$ with $m_{0}\in\Phi^{-1}(0)$ and $g\in G^{\mathbb{C}}$. By the Cartan decomposition of $G^{\mathbb{C}}$, we have $G^{\mathbb{C}}=PG$, where $G$ is a maximal compact subgroup of $G^{\mathbb{C}}$ and the Lie algebra of $P$ is $\sqrt{-1}\mathfrak{g}$. Thus we write $g={\rm e}^{\eta}k$ with $\eta=\sqrt{-1}\xi\in\sqrt{-1}\mathfrak{g}$ and $k\in G$. Replacing $m_{0}$ with $km_{0}$, we can assume that $m={\rm e}^{\eta}m_{0}$.
Consider the curve $\gamma\colon(-\infty,+\infty)\rightarrow M$, $t\mapsto\gamma(t)={\rm e}^{t\eta}m_{0}$. Hence, by the formula $\eta_{M}\langle s,s\rangle=-4\pi\langle\Phi(\cdot),\xi\rangle\langle s,s\rangle$ we have, along $\gamma(t)$, $\frac{{\rm d}}{{\rm d}t}\langle s,s\rangle=-4\pi\langle\Phi(\cdot),\xi\rangle\langle s,s\rangle$. As in [9, Lemma 4.7], we note that $\eta_{M}$ is the gradient vector field associated with the function ${\Phi^{\xi}:=\langle\Phi(\cdot),\xi\rangle}$. Thus $\Phi^{\xi}$ is strictly increasing along $\gamma(t)$, so it is positive for $t>0$ and negative for $t<0$. Therefore, $\langle s,s\rangle$ has a unique maximum mum at $t=0$. ∎
Finally, we note that Theorem 5.10 implies the surjectivity of $\mathfrak{r}$. Indeed, if $m$ is a point of $M$, then we can find a neighborhood $U$ of $m$ in $M$ and a non-zero holomorphic $G$-invariant $L^{k}$-valued $(0,n_{-})$-forms, $s_{0}\colon U\rightarrow L^{k}\otimes T^{*0,n_{-}}M$. Then $s=fs_{0}$ on $U\cap M_{s}$, $f$ being a bounded holomorphic function. Since $(M\setminus M_{s})\cap U$ is of complex codimension greater than one in $U$, the singularity of $f$ at $m$ is removable. Thus $s$ extends to a holomorphic section of $L^{k}\otimes T^{*0,n_{-}}M$ over all of $M$.
@@
$\big{\lVert}I_{k}^{G}[f]\big{\rVert}_{\infty}\leq\big{\rVert}T^{G}_{k}\big{[}\widetilde{f}\big{]}\big{\lVert}\leq\lVert f\rVert_{\infty}.$ (6.1)
Recall that given two $G$ invariant $(0,n_{-})$ forms $s$ and $t$ in $H(X)^{G}_{k}$, we denote their inner product by $(s|t)$. Now, using the Cauchy–Schwartz inequality, we obtain that for each element ${x\in\mu^{-1}(0)}$
$\big{\lVert}I^{G}_{k}[f](x)\big{\rVert}^{2}_{\infty}$ $=\frac{\big{|}\bigl{(}S^{G}_{k}(\cdot,x)v\big{|}T^{G}_{k}[f]S^{G}_{k}(\cdot,x)v\bigr{)}\big{|}^{2}}{\bigl{(}S^{G}_{k}(\cdot,x)v\big{|}S^{G}_{k}(\cdot,x)v\bigr{)}^{2}}$
$\leq\frac{\bigl{(}T^{G}_{k}[f]S^{G}_{k}(\cdot,x)v\big{|}T^{G}_{k}[f]S^{G}_{k}(\cdot,x)v\bigr{)}}{\bigl{(}S^{G}_{k}(\cdot,x)v\big{|}S^{G}_{k}(\cdot,x)v\bigr{)}}\leq\rVert T^{G}_{k}[f]\lVert^{2},$
where the last inequality follows by the definition of the operator norm, see (2.3). Taking the $\sup$ both sides over $\mu^{-1}(0)$, it shows the first inequality in (6.1).
@@
$\frac{\lVert T^{G}_{k}[f]s\rVert^{2}}{\lVert s\rVert^{2}}$ $=\frac{\bigl{(}T^{G}_{k}[f]s\big{|}T^{G}_{k}[f]s\bigr{)}}{\int_{X}\langle s,s\rangle\mathrm{dV}_{X}}=\frac{\lVert f\rVert^{2}_{\infty}\bigl{(}S^{G}_{k}[f]s\big{|}S^{G}_{k}[f]s\bigr{)}}{\int_{X}\langle s,s\rangle\mathrm{dV}_{X}}\leq\lVert f\rVert^{2}_{\infty}.$
Hence, it proves the second inequality (6.1).
Now, we are ready to prove the theorem. Since $X$ is compact, then $\mu^{-1}(0)$ is closed and hence compact. Choose $x_{e}\in\mu^{-1}(0)$ a point with $\big{\lvert}\widetilde{f}(x_{e})\big{\rvert}=\lVert f\rVert_{\infty}$. Since the Berezin transform has as a leading term the identity, we have $\big{\lvert}\bigl{(}I^{G}_{k}[f]\bigr{)}(x_{e})-\widetilde{f}(x_{e})\big{\rvert}\leq\frac{C}{k}$ for a suitable $C>0$. Thus,
$\lVert f\rVert_{\infty}-\frac{C}{k}=\big{\lvert}\widetilde{f}(x_{e})\big{\rvert}-\frac{C}{k}\leq\big{\lvert}\bigl{(}I^{G}_{k}[f]\bigr{)}(x_{e})\big{\rvert}\leq\big{\lVert}I^{G}_{k}[f]\big{\rVert}_{\infty}.$
Putting (6.1) in this last inequality, we get part (a) of Theorem 2.3.
Parts (b) and (c) of Theorem 2.3 are consequence of Theorem 3.5.
7 Proof of Theorem 1.3
It is an easy consequence of [7] that the star product defined by Toeplitz operators is of Wick type.
Let $f$ and $g$ be two $G$ and $S^{1}$ invariant smooth function on $X_{G}$. For ease of notation, we write $T^{G}_{k}[f]$ for $T^{G}_{k}\big{[}\widetilde{f}\big{]}$. It is clear by Theorem 2.3 and [7] that $\star$ is a well-defined associative star product. In fact, by part (b) of Theorem 2.3, we get
$\big{\lVert}T^{G}_{k}[{\{f,g\}-i(C_{1}(f,g)-C_{1}(g,f))}]\big{\rVert}=O\left(\frac{1}{k}\right),$
and taking the limit for $k\rightarrow+\infty$ and using part (a) we get $\lVert\{f,g\}-{\rm i}(C_{1}(f,g)-C_{1}(g,f))\rVert_{\infty}=0$ and thus we get part (b). Similarly, by Theorem 2.3 (c) and (a), we get that $C_{0}(g,h)=g\cdot h$. In a similar way the associativity property is proved.
Now, we prove the uniqueness. Let $C_{j}$ and $\tilde{C}_{j}$ be two systems of bi-differential operators inducing star products and both fulfilling the asymptotic condition (1.1). We show that $C_{j}=\tilde{C}_{j}$, for all $j\in\mathbb{N}_{0}$, hence the induced star products coincide. Note $C_{0}=\tilde{C}_{0}$ by hypothesis. By induction, assume that $C_{j}=\tilde{C}_{j}$ for each $j\leq N-2$. By the results in [7], we have
$T_{k}^{G}[f]\circ T_{k}^{G}[g]\sim\sum_{j=0}^{N-1}k^{-j}T_{k}^{G}[C_{j}(f,g)]+O\bigl{(}k^{-N}\bigr{)}$
@@
$T_{k}^{G}[f]\circ T_{k}^{G}[g]\sim\sum_{j=0}^{N-1}k^{-j}T_{k}^{G}\big{[}\tilde{C}_{j}(f,g)\big{]}+O\bigl{(}k^{-N}\bigr{)},$
@@
$\left\lVert\frac{1}{k^{N-1}}T^{G}_{k}\big{[}{C}_{N-1}(f,g)-\tilde{C}_{N-1}(f,g)\big{]}\right\rVert\leq\frac{K}{k^{N}}.$
Eventually, taking the limit on both sides as $k$ goes to infinity and applying Theorem 2.3 (a), we obtain $C_{N-1}(f,g)=\tilde{C}_{N-1}(f,g)$.
Now we recall that
$A_{N}=\hat{R}^{l}T^{G}[f]T^{G}[g]-\sum_{j=0}^{N-1}\hat{R}^{N-j}T^{G}[C_{j}(f,g)]$
which is an operator of Szegő type of order zero by Theorem 3.4 and it is invariant under the circle action. Thus, its symbol is a $G$ and $S^{1}$ invariant function on $X_{G}$ which, by restriction on $\mu^{-1}(0)$, defines a $S^{1}$-invariant function on $X_{G}$. By definition, it is the next element $C_{N}(f,g)$ in the star-product.
@@
$C_{j}(1,f)=C_{j}(f,1)=0$ (7.1)
for $j\geq 1$ and for every $f\in C^{\infty}(X_{G})^{S^{1}}$. First, by Theorem 3.2 note that $C_{0}(1,g)=C_{0}(1,g)=g$ for every $k$. Now, if we put $g=1$,
$A_{1}=\hat{R}T^{G}[f]T^{G}[g]-\hat{R}T^{G}[f\cdot g]=0,$
and hence the symbol of $A_{1}$ vanishes. By induction, the claim (7.1) follows.
Let us now prove that the Berezin–Toeplitz star product fulfills parity. Considering the formal parameter to be real, $\nu=\overline{\nu}$, this is equivalent to
$\overline{C_{k}(f,g)}=C_{k}(\overline{g},\overline{f})$ (7.2)
for $k\geq 0$. Note that, as an easy consequence of the definition of Berezin–Toeplitz star product, we have $T_{k}^{G}[f]^{\dagger}=T_{k}^{G}[\overline{f}]$. The star product $\overline{g}\star\overline{f}$ is given by the asymptotic expansion of the composition
$T_{k}^{G}[\overline{g}]\cdot T_{k}^{G}\big{[}\overline{f}\big{]}$ $=T_{k}^{G}[g]^{{\dagger}}\cdot T_{k}^{G}[f]^{{\dagger}}=\bigl{(}T_{k}^{G}[{f}]\cdot T_{k}^{G}[g]\bigr{)}^{{\dagger}}$
$\sim\sum_{j=0}^{+\infty}k^{-j}T_{k}^{G}[C_{j}(f,g)]^{{\dagger}}\sim\sum_{j=0}^{+\infty}k^{-j}T_{k}^{G}\big{[}\overline{C_{j}(f,g)}\big{]}.$
Thus, the claim (7.2) follows.
By [7], see also [8], the trace $\mathrm{Tr}_{k}$ on $H(X)^{G}_{k}$ admits a complete asymptotic expansion in decreasing integer power of $k$
$\mathrm{Tr}_{k}\bigl{(}T^{G}_{k}[f]\bigr{)}\sim k^{d}\left(\sum_{k=0}^{+\infty}k^{-j}\tau_{j}(f)\right),$
with $\tau_{j}(f)\in\mathbb{C}$. It induces a $\mathbb{C}[[\nu]]$-linear map $\mathrm{Tr}\colon{C}^{\infty}(X_{G})^{S^{1}}[[\nu]]\rightarrow\nu^{-n}\mathbb{C}[[\nu]]$ such that
@@
$\mathrm{Tr}(f\star g)=\mathrm{Tr}(g\star f).$ (7.3)
By $\mathbb{C}[[\nu]]$-linearity, we prove (7.3) for $f$ and $g$ in ${C}^{\infty}(X_{G})^{S^{1}}$. Note that $f\star g-g\star f$ is given by the asymptotic expansion of $T^{G}_{k}[f]T^{G}_{k}[g]-T^{G}_{k}[g]T^{G}_{k}[f]$. Hence the trace of $f\star g-g\star f$ is given by the expansion of $\mathrm{Tr}_{k}\bigl{(}T^{G}_{k}[f]T^{G}_{k}[g]-T^{G}_{k}[g]T^{G}_{k}[f]\bigr{)}$. But for every $k$ this vanishes and thus we get (7.3).
Acknowledgements
We are indebted to the referees for various interesting comments and for suggesting several improvements. The author thanks Chin-Yu Hsiao and Herrmann Hendrik for their valuable conversations. The author is a member of GNSAGA, part of the Istituto Nazionale di Alta Matematica, and expresses gratitude to the group for its support.